# Angular bispectrum of matter number counts in cosmic structures

### v2

@@ 6 Conclusions
In this article, starting from the first- and second-order number count results (Yoo et al., 2009; Yoo, 2010; Challinor and Lewis, 2011; Bonvin and Durrer, 2011; Jeong et al., 2012; Yoo and Zaldarriaga, 2014; Bertacca et al., 2014; Di Dio et al., 2014; Magi and Yoo, 2022), we have numerically evaluated the redshift binned angular bispectrum of the galaxy number count without relying on the Limber approximation. Our calculation includes redshift binning and consistently incorporates all Newtonian terms (density, redshift-space distortions, and quadratic terms), non-integrated projection effects, as well as all dynamical relativistic and radiative contributions. By extending the computation of Assassi et al. (2017), we reduced the problem to one-dimensional integrals, making the numerical computation of the considered bispectrum contributions feasible. In total, we have incorporated 7 linear and 14 second-order terms (summarised in Eqs. 4 and 2.2 respectively) and evaluated the resulting bispectrum for various triangle configurations. Our main analytical bispectrum results – presented in a computationally tractable form – are summarised in Eq. (3), with the individual contributions given in Eqs. (3.1)–(42).
@@ Appendix B Second-order perturbation theory kernels
In second-order perturbation theory, fields can be written in Fourier space as a convolution integral such as Eqs. (32) and (33). Hence we can write
$\mathcal{I}_{2}^{\rm X}(r_{1},\boldsymbol{k}_{1})$ $=\mathcal{N}^{2}D^{2}(r_{1})\int\frac{d^{3}k_{2}d^{3}k_{3}}{(2\pi)^{3}}\delta_{\mathrm{D}}^{(3)}(\boldsymbol{k}_{1}-\boldsymbol{k}_{2}-\boldsymbol{k}_{3})k_{2}^{2}k_{3}^{2}\mathcal{K}_{\rm X}^{\mathcal{I}}(r_{1},k_{1},k_{2},k_{3})\phi_{0}(\boldsymbol{k}_{2})\phi_{0}(\boldsymbol{k}_{3})\,$, (58)
where the index $\rm X$ stands for Newtonian (N), general relativistic (GR) or radiation (Rad) and $\mathcal{I}$ can be $\delta$ or $v$. In this section, we give the expressions of the kernels that are needed for our derivation. Following Tram et al. (2016), the kernels defined in Eqs. (32) and (33) read (Villa and Rampf, 2016; Tram et al., 2016)
@@
$\alpha^{\delta}_{\rm Rad}=0,\quad\beta^{\delta}_{\rm Rad}=0,\quad\gamma^{\delta}_{\rm Rad}=-\frac{1}{2}\left(f+\frac{3\Omega_{m}}{2}\right)\left(\frac{\mathcal{H}^{2}}{k_{1}^{2}}+3f\frac{\mathcal{H}^{4}}{k_{1}^{4}}\right)\frac{\partial\log{T_{\phi}}}{\partial\log{k_{1}}}\,.$ (62)
@@
$-\frac{8D\dot{D}}{aH_{0}^{2}\Omega_{m0}}\Theta_{0}$ $\longmapsto k_{1}^{2}k_{2}^{2}\mathcal{N}^{2}D^{2}9f\Omega_{m}\frac{\mathcal{H}^{3}}{k^{4}}\left[1+\frac{1}{3}\mu^{2}+\frac{2}{3}\left(\frac{k_{1}}{k_{2}}+\frac{k_{2}}{k_{1}}\right)\mu\right]$
@@
$\frac{8\dot{F}}{9(H_{0}^{2}\Omega_{m0})^{2}}\Psi_{0}$ $\longmapsto k_{1}^{2}k_{2}^{2}\mathcal{N}^{2}D^{2}\frac{6w}{7}\frac{\mathcal{H}}{k^{2}}\left[1-\mu^{2}\right]$ (63)
@@
$\frac{f^{(-2)}_{00}}{\mathcal{H}^{2}}$ $=\frac{\beta_{1}-\alpha_{1}}{2}+2\gamma_{1}\,$, $\frac{f^{(-2)}_{2,-2}}{\mathcal{H}^{2}}$ $=\frac{\alpha_{1}-\beta_{1}}{4}+3\gamma_{1}\,$, $f^{(0)}_{0,-2}$ $=\frac{\mathcal{H}^{2}}{2}\left(\frac{\beta_{1}}{2}-\alpha_{1}\right)\,$,
@@
For the radiation part $\gamma_{\rm Rad}$, we can use the same kind of decomposition for the factor in front of $\partial\log T_{\phi}/\partial\log k_{1}$: $\gamma^{R}=\sum_{i}\gamma_{i}^{R}(\mathcal{H}/k)^{2i}$. We use FFTLog to decompose $\partial\log T_{\phi}/\partial\log k_{1}$ such that
$f_{00}^{(-2,R)}=-2\gamma_{1}^{R}\mathcal{H}^{2}\,$$\qquad$ $f_{2,-2}^{(-2,R)}=\gamma_{1}^{R}\mathcal{H}^{2}\,$,
$f_{00}^{(-4,R)}=-2\gamma_{2}^{R}\mathcal{H}^{4}\,$$\qquad$ $f_{2,-2}^{(-4,R)}=\gamma_{2}^{R}\mathcal{H}^{4}\,.$ (68)
@@ Appendix D Numerical implementation
  • Now we precompute the generalised power spectra $C_{\ell}(\chi)$ defined in Eqs. (30). Note that in practice, we only need a few generalised power spectra to /obtain all of them. In total, we compute $7$ spectra for all $\ell$ and for a chosen number of $\chi$. This part is the bottleneck of the whole computation.

### v3

@@ 6 Conclusions
In this article, starting from the first- and second-order number count results (Yoo et al., 2009; Yoo, 2010; Challinor and Lewis, 2011; Bonvin and Durrer, 2011; Jeong et al., 2012; Yoo and Zaldarriaga, 2014; Bertacca et al., 2014; Di Dio et al., 2014; Magi and Yoo, 2022), we have numerically evaluated the redshift binned angular bispectrum of the galaxy number count without relying on the Limber approximation. Our calculation includes redshift binning and consistently incorporates all Newtonian terms (density, redshift-space distortions, and quadratic terms), non-integrated projection effects, as well as all dynamical relativistic and radiative contributions. Integrated terms, such as lensing, can be neglected for auto-bispectra (Di Dio et al., 2016a), and removed from our simulations measurements. By extending the computation of Assassi et al. (2017), we reduced the problem to one-dimensional integrals, making the numerical computation of the considered bispectrum contributions feasible. In total, we have incorporated 7 linear and 14 second-order terms (summarised in Eqs. 4 and 2.2 respectively) and evaluated the resulting bispectrum for various triangle configurations. Our main analytical bispectrum results – presented in a computationally tractable form – are summarised in Eq. (3), with the individual contributions given in Eqs. (3.1)–(42).
@@ Appendix B Second-order perturbation theory kernels
In second-order perturbation theory, fields can be written in Fourier space as a convolution integral such as Eqs. (32) and (33) where the kernel can be denotted in general as $\mathcal{K}_{\rm X}^{\mathcal{I}}$, where the index $\rm X$ stands for Newtonian (N), general relativistic (GR) or radiation (Rad), and $\mathcal{I}$ can be $\delta$ or $v$. In this section, we give the expressions of the kernels that are needed for our derivation. Following Tram et al. (2016), the kernels defined in Eqs. (32) and (33) read (Villa and Rampf, 2016; Tram et al., 2016)
@@
$\alpha^{\delta}_{\rm Rad}=0,\quad\beta^{\delta}_{\rm Rad}=0,\quad\gamma^{\delta}_{\rm Rad}=-\frac{1}{2}\left(f+\frac{3\Omega_{m}}{2}\right)\left(\frac{\mathcal{H}^{2}}{k_{1}^{2}}+3f\frac{\mathcal{H}^{4}}{k_{1}^{4}}\right)\frac{\partial\log{T_{\phi}}}{\partial\log{k_{1}}}\,$, (61)
where $T_{\phi}$ stands for the transfer function of the gravitational potential $\phi$.
@@
$\frac{4D\dot{D}}{aH_{0}^{2}\Omega_{m0}}\Delta^{-1}\left[\Delta^{-1}\left(\phi_{0}^{l}\phi_{0}^{m}\right)_{,lm}-\frac{1}{3}\phi_{0}^{l}\phi_{0,l}\right]$ $\longmapsto k_{1}^{2}k_{2}^{2}\mathcal{N}^{2}D^{2}9f\Omega_{m}\frac{\mathcal{H}^{3}}{k^{4}}\left[1+\frac{1}{3}\mu^{2}+\frac{2}{3}\left(\frac{k_{1}}{k_{2}}+\frac{k_{2}}{k_{1}}\right)\mu\right]$
@@
$\frac{4\dot{F}}{9(H_{0}^{2}\Omega_{m0})^{2}}\Delta^{-1}\left[\phi_{0,lm}\phi_{0}^{lm}-(\Delta\phi_{0})^{2}\right]$ $\longmapsto k_{1}^{2}k_{2}^{2}\mathcal{N}^{2}D^{2}\frac{6w}{7}\frac{\mathcal{H}}{k^{2}}\left[1-\mu^{2}\right]$ (62)
@@
$\frac{f^{(-2)}_{00}}{\mathcal{H}^{2}}$ $=\frac{\beta_{1}-\alpha_{1}}{2}-2\gamma_{1}\,$, $\frac{f^{(-2)}_{2,-2}}{\mathcal{H}^{2}}$ $=\frac{\alpha_{1}-\beta_{1}}{4}+\gamma_{1}\,$, $f^{(0)}_{0,-2}$ $=\frac{\mathcal{H}^{2}}{2}\left(\frac{\beta_{1}}{2}-\alpha_{1}\right)\,$,
@@
For the radiation part $\gamma_{\rm Rad}$, we can use the same kind of decomposition for the factor in front of $\partial\log T_{\phi}/\partial\log k_{1}$: $\gamma_{\rm Rad}=\sum_{i}\gamma_{i}^{\rm Rad}(\mathcal{H}/k)^{2i}$. We use FFTLog to decompose $\partial\log T_{\phi}/\partial\log k_{1}$ such that
$f_{00}^{(-2,R)}=-2\gamma^{\rm Rad}_{1}\mathcal{H}^{2}\,$$\qquad$ $f_{2,-2}^{(-2,R)}=\gamma^{\rm Rad}_{1}\mathcal{H}^{2}\,$,
$f_{00}^{(-4,R)}=-2^{\rm Rad}_{2}\mathcal{H}^{4}\,$$\qquad$ $f_{2,-2}^{(-4,R)}=\gamma^{\rm Rad}_{2}\mathcal{H}^{4}\,.$ (67)
@@ Appendix D Numerical implementation
  • Now we precompute the generalised power spectra $C_{\ell}(\chi)$ defined in Eqs. (30). Note that in practice, we only need a few generalised power spectra to obtain all of them. In total, we compute $7$ spectra for all $\ell$ and for a chosen number of $\chi$. This part is the bottleneck of the whole computation.